# MiMo-V2-Flash Technical Report

### v1

@@ 1 Introduction
Recent progress towards Artificial General Intelligence (AGI) is increasingly propelled by two frontiers: advanced reasoning chains and autonomous agentic workflows (gemini3pro; liu2025deepseek; team2025kimi), grounded in large-scale Reinforcement Learning (RL). Yet building scalable reasoners and agents hits a common critical bottleneck, where long-context modeling must be simultaneously fast and strong.
In this work, we introduce MiMo-V2-Flash, an efficient and cost-effective Large Language Model (LLM) that delivers strong reasoning and agentic performance. MiMo-V2-Flash is a 309B-parameter MoE with 15B activated per token. To alleviate the quadratic complexity of full attention, MiMo-V2-Flash adopts a hybrid attention mechanism that interleaves local sliding window and global attention. The sliding window size is 128-token and the hybrid local:global ratio is 5:1, yielding nearly a 6× reduction in KV-cache storage and attention computation for long contexts. With the help of learnable attention sink bias (agarwal2025gpt), the hybrid architecture maintains strong modeling capability even in long-context scenarios, despite the aggressive sliding window size and hybrid ratio. MiMo-V2-Flash also incorporates Multi-Token Prediction (MTP) to enhance training performance and accelerate inference decoding. In particular, MTP has strong potential to boost RL rollout speed, which helps to scale LLMs towards greater intelligence. With a lightweight dense Feed-Forward Network (FFN) and sliding window attention, our MTP block delivers substantial decoding speedups in practice at high acceptance rates.
The pre-training recipe of MiMo-V2-Flash largely follows that of MiMo-7B (xia2025mimo), with several enhancements. Training is conducted using FP8 mixed-precision, enabling efficient large-scale training over 27T tokens. The model is initially pre-trained with a native 32K context and later extended to 256K. The resulting pretrained model, MiMo-V2-Flash-Base, has been evaluated against leading open-source base models such as Kimi-K2-Base (team2025kimik2) and DeepSeek-V3.2-Exp-Base (liu2025deepseek). MiMo-V2-Flash-Base achieves competitive performance across general benchmarks and surpasses peer models on reasoning-focused tasks. For long-context retrieval, our hybrid attention architecture achieves nearly 100% success rates across context lengths from 32K to 256K. On the extreme long-context reasoning benchmark GSM-Infinite (zhou2025gsm), MiMo-V2-Flash demonstrates robust performance with minimal degradation when scaling from 16K to 128K.
@@ 2.1 Overall Architecture
As illustrated in Figure 2, MiMo-V2-Flash follows a standard Transformer (vaswani2017attention) backbone augmented with MoE (shazeer2017outrageously) and hybrid attention (gpt3; gemma2; team2025gemma; li2025minimax; qwen3next2025; kda). MiMo-V2-Flash is mainly composed of repeated hybrid blocks that interleave Local Sliding Window Attention (SWA) and Global Attention (GA). It stacks $M=8$ hybrid blocks, each structured with $N=5$ consecutive SWA blocks followed by an GA block. The only exception is the very first Transformer block, which uses global attention with a dense Feed-Forward Network (FFN) to stabilize early representation learning. The sliding window size $W$ used in MiMo-V2-Flash is $128$. Both the SWA block and the GA block utilize the sparse MoE FFN. Each MoE layer comprises 256 experts in total, with 8 activated per token, and contains no shared experts.
MiMo-V2-Flash also integrates MTP (gloeckle2024better; liu2024deepseek; xia2025mimo) to improve model performance (both quality and efficiency). Worth noting, the MTP block uses dense FFN instead of MoE and applies SWA rather than GA, making it lightweight for speculative decoding. The number of parameters for each MTP block is only 0.33B.
@@
Table 1 summarizes detailed configurations of MiMo-V2-Flash. The model consists of 39 SWA layers and 9 GA layers. Both SWA and GA utilize Grouped-Query Attention (GQA) (ainslie2023gqa). Specifically, SWA has 64 query heads and 8 key-value heads, while GA has 64 query heads and 4 key-value heads. The per-head dimensions are the same for SWA and GA (192 for queries and keys, and 128 for values). Rotary Positional Embedding (RoPE, su2024roformer) is partially applied to the first 64 dimensions query and key. Following recent best practices, we adopt an FP8 mixed-precision framework similar to DeepSeek-V3 (liu2024deepseek). Specifically, we retain BF16 precision for the attention output projections, as well as for the embedding and output head parameters, while maintaining FP32 precision for the MoE router parameters. This mixed-precision configuration improves numerical stability without materially impacting training efficiency or memory footprint.
@@ 2.2 Hybrid Sliding Window Attention Architecture
Sliding window attention (beltagy2020longformer) restricts each token’s attention scope to a local window rather than the entire sequence, thereby reducing both computational and memory complexity dramatically. This naturally motivates hybrid attention architectures that interleave sliding window attention with global attention. However, prior work has shown that overly aggressive use of SWA, such as very small sliding window sizes or high SWA:GA ratios, can lead to substantial degradation in model performance (team2025gemma), especially in long-context tasks. Recently, the introduction of learnable attention sink bias, which allows the model to assign little or no attention to tokens when needed, has substantially enhanced the modeling capacity of SWA-based architectures (agarwal2025gpt). While the precise theoretical underpinnings of the attention sink mechanism remain an active research area (xiao2023efficient; attnsink_study; qiu2025gatedattn; sun2024massive), we empirically observed that learnable attention sinks bias dramatically enhance the performance of hybrid SWA models, matching or even surpassing baselines with fully GA layers.
In MiMo-V2-Flash, our implementation follows the design used in gpt-oss (agarwal2025gpt), where a learnable attention sink bias $sink\in\mathbb{R}$ is applied to the denominator of softmax for each attention head. Specifically, let the attention logits between token $i$ and $j$ of one single head be:
@@ 2.2.1 Model Architecture Experiments
We evaluate four model architecture variants in a comparative setting. These include an all global attention (All GA) baseline, a hybrid SWA model with a 128-token window without attention sinks bias, and two hybrid SWA models augmented with attention sinks bias using window sizes of 128 and 512, respectively. All variants share the same training pipeline: pre-training on 250B tokens with an 8,192 sequence length, long context extension to 32,768 over an additional 40B tokens, followed by long-context SFT and reasoning SFT with chain-of-thought supervision. We evaluate model variants across benchmarks covering general capability, long-context understanding, and complex reasoning. General-domain results (Table 2) are obtained from pre-trained base models without long-context extension, evaluating general knowledge and reasoning on MMLU hendrycks2020measuring, BBH (suzgun2022challenging), TriviaQA (joshi2017triviaqa), GSM8K (cobbe2021training), MATH (hendrycks2021measuring), CMMLU (li2023cmmlu), and MBPP (austin2021program). Long-context results (Table 3) evaluate long-context–extended base models on GSM-Infinite (zhou2025gsm), NoLiMa (modarressi2025nolima), and RULER-32k (hsieh2024ruler), and long-context SFT models on MRCR (vodrahalli2024michelangelo). For GSM-Infinite and Nolima, we construct internal few-shot benchmarks to assess base models under controlled long-context settings. Complex reasoning results (Table 4) evaluate the reasoning SFT models on AIME24&25 (AIME), LiveCodeBench (jain2024livecodebench), and GPQA-Diamond (rein2024gpqa).
@@ 2.3.1 Motivation of using MTP
Prior work demonstrates that MTP is a powerful training objective that enhances training efficiency and model quality (gloeckle2024better; liu2024deepseek; xia2025mimo). Beyond these training benefits, we place stronger emphasis on exploiting MTP as a native draft model for self-speculative decoding to deliver real-deployment speedup. In the following, we elaborate on how MTP accelerates inference from two perspectives: general LLM decoding speedup and RL training acceleration.
@@
MTP acceleration is particularly well-suited for RL training (miles2025), where the rollout phase consistently emerges as the dominant bottleneck due to the inference and decoding costs. MTP addresses two key challenges in RL training:
  • It enables efficient and effective RL with small batches. Current RL training relies on large-batch, off-policy algorithms to maximize throughput (schulman2017proximal; zheng2025group; liu2025deepseek). However, on-policy training is generally more stable and effective, yet its small batches underutilize GPU resources. MTP mitigates this limitation by scaling token-level parallelism instead of batch size, making small-batch, on-policy RL training more practical.
  • It mitigates GPU idleness from long-tail stragglers. As the rollout phase progresses, long-tail stragglers that process long sequences with small batch sizes (often approaching 1) can cause significant GPU idleness (zhong2025optimizing; gao2025rollpacker). In such scenarios, MTP enhances the computational efficiency of both attention and FFN, substantially reducing overall latency.
@@ 3 Pre-Training
The MiMo-V2-Flash pre-training corpus consists of 27 trillion tokens drawn from a diverse collection of high-quality sources, including public web content, books, academic papers, code, mathematics, and broader STEM materials. Our data processing pipeline largely follows that of MiMo-7B (xia2025mimo), with a deliberate shift toward data exhibiting long-range dependencies. In particular, we emphasize long-form web documents and carefully curated code corpora such as repository-level code, pull requests, issues, and commit histories to strengthen the model’s ability to capture extended contextual relationships and perform complex, multi-step reasoning.
@@ 3.3.1 Evaluation Setup
We evaluate MiMo-V2-Flash-Base on a series of benchmarks, encompassing various capabilities: (1) General language understanding and reasoning, including BBH (suzgun2022challenging), MMLU (hendrycks2020measuring), MMLU-Redux (gema2024we), MMLU-Pro (wang2024mmlu), DROP (dua2019drop), ARC (clark2018think), HellaSwag (zellers2019hellaswag), WinoGrande (sakaguchi2021winogrande), TriviaQA (joshi2017triviaqa), GPQA-Diamond (rein2024gpqa), SuperGPQA (du2025supergpqa), and SimpleQA (simpleqa). (2) Mathematics reasoning. GSM8K (cobbe2021training), MATH (hendrycks2021measuring), and AIME (AIME) (2024 & 2025). (3) Coding. HumanEval+ (liu2023your), MBPP+ (liu2023your), CRUXEval (gu2024cruxeval), MultiPL-E (cassano2022multipl), BigCodeBench (zhuo2024bigcodebench), LiveCodeBench-v6 (jain2024livecodebench), and SWE-Bench (jimenez2023swe) (few-shot Agentless Repair (xia2024agentless)). (4) Chinese understanding. C-Eval (huang2023c), CMMLU (li2023cmmlu), and C-SimpleQA (he2025chinese). (5) Multilingual understanding. GlobalMMLU (singh2025global), and INCLUDE (romanou2024include). (6) Long context. NIAH-Multi (hsieh2024ruler), GSM-Infinite (zhou2025gsm) (5-shot, Hard Ops-{2,4,6,8,10}).
@@ 3.3.2 Evaluation Results
Table 5 presents a comprehensive comparison of MiMo-V2-Flash-Base against leading open-source base models (team2025kimik2; liu2024deepseek). MiMo-V2-Flash-Base delivers competitive performance across most benchmarks and consistently outperforms peers on reasoning tasks (MMLU-Pro, GPQA-Diamond, AIME). On SWE-Bench, it even surpasses the substantially larger Kimi-K2-Base while using less than one-third the parameters, underscoring the strength of our approach for realistic code-agent tasks. However, constrained by its limited parameter count, we observe MiMo-V2-Flash exhibits lower knowledge capacity compared to larger models, as reflected in SimpleQA.
@@ 4.1 Multi-Teacher On-Policy Distillation (MOPD): A New Post-Training Paradigm
Rather than merging model parameters or generating static offline datasets from experts, we formulate multi-teacher knowledge integration as an on-policy reinforcement learning process. The student model samples from its own evolving distribution and receives token-level supervision from domain-specific teachers through KL divergence rewards (minillm; Agarwal2023OnPolicyDO; lu2025onpolicydistillation), effectively combining specialized capabilities without the traditional trade-offs (Table 7).
@@ 4.4 Technical Formulation of MOPD
By default, we combine the advantages of MOPD with other types of advantages, such as those computed using Outcome Reward Models (ORMs), including GRPO (shao2024deepseekmath). Let $\hat{A}_{\text{ORM}}$ denote the advantages computed by the ORMs; the final advantages are given by:
@@ 4.5.1 Evaluation Setup
We evaluate MiMo-V2-Flash on MMLU-Pro (wang2024mmlu), GPQA-Diamond (rein2024gpqa), HLE Text-only (phan2025humanity), AIME 2025 (AIME), LiveCodeBench (2024.08-2025.04) (jain2024livecodebench), HMMT Feb. 2025 (balunovic2025matharena), Arena-Hard (arenahard2024), LongBench V2 (bai2025longbench), MRCR (vodrahalli2024michelangelo) ({2,4,8}-needles, maximum 128K), SWE-Bench Verified (jimenez2024swebench), SWE-Bench Multilingual (yang2025swesmith), Terminal-Bench, BrowseComp (wei2025browsecomp), $\tau^{2}$-Bench (barres2025tau).

### v2

@@ 1 Introduction
Recent progress towards Artificial General Intelligence (AGI) is increasingly propelled by two frontiers: advanced reasoning chains and autonomous agentic workflows [gemini3pro, liu2025deepseek, team2025kimi], grounded in large-scale Reinforcement Learning (RL). Yet building scalable reasoners and agents hits a common critical bottleneck, where long-context modeling must be simultaneously fast and strong.
In this work, we introduce MiMo-V2-Flash, an efficient and cost-effective Large Language Model (LLM) that delivers strong reasoning and agentic performance. MiMo-V2-Flash is a 309B-parameter MoE with 15B activated per token. To alleviate the quadratic complexity of full attention, MiMo-V2-Flash adopts a hybrid attention mechanism that interleaves local sliding window and global attention. The sliding window size is 128-token and the hybrid local:global ratio is 5:1, yielding nearly a 6× reduction in KV-cache storage and attention computation for long contexts. With the help of learnable attention sink bias [agarwal2025gpt], the hybrid architecture maintains strong modeling capability even in long-context scenarios, despite the aggressive sliding window size and hybrid ratio. MiMo-V2-Flash also incorporates Multi-Token Prediction (MTP) to enhance training performance and accelerate inference decoding. In particular, MTP has strong potential to boost RL rollout speed, which helps to scale LLMs towards greater intelligence. With a lightweight dense Feed-Forward Network (FFN) and sliding window attention, our MTP block delivers substantial decoding speedups in practice at high acceptance rates.
The pre-training recipe of MiMo-V2-Flash largely follows that of MiMo-7B [xia2025mimo], with several enhancements. Training is conducted using FP8 mixed-precision, enabling efficient large-scale training over 27T tokens. The model is initially pre-trained with a native 32K context and later extended to 256K. The resulting pretrained model, MiMo-V2-Flash-Base, has been evaluated against leading open-source base models such as Kimi-K2-Base [team2025kimik2] and DeepSeek-V3.2-Exp-Base [liu2025deepseek]. MiMo-V2-Flash-Base achieves competitive performance across general benchmarks and surpasses peer models on reasoning-focused tasks. For long-context retrieval, our hybrid attention architecture achieves nearly 100% success rates across context lengths from 32K to 256K. On the extreme long-context reasoning benchmark GSM-Infinite [zhou2025gsm], MiMo-V2-Flash demonstrates robust performance with minimal degradation when scaling from 16K to 128K.
@@ 2.1 Overall Architecture
As illustrated in Figure 2, MiMo-V2-Flash follows a standard Transformer [vaswani2017attention] backbone augmented with MoE [shazeer2017outrageously] and hybrid attention [gpt3, gemma2, team2025gemma, li2025minimax, qwen3next2025, kda]. MiMo-V2-Flash is mainly composed of repeated hybrid blocks that interleave Local Sliding Window Attention (SWA) and Global Attention (GA). It stacks $M=8$ hybrid blocks, each structured with $N=5$ consecutive SWA blocks followed by an GA block. The only exception is the very first Transformer block, which uses global attention with a dense Feed-Forward Network (FFN) to stabilize early representation learning. The sliding window size $W$ used in MiMo-V2-Flash is $128$. Both the SWA block and the GA block utilize the sparse MoE FFN. Each MoE layer comprises 256 experts in total, with 8 activated per token, and contains no shared experts.
MiMo-V2-Flash also integrates MTP [gloeckle2024better, liu2024deepseek, xia2025mimo] to improve model performance (both quality and efficiency). Worth noting, the MTP block uses dense FFN instead of MoE and applies SWA rather than GA, making it lightweight for speculative decoding. The number of parameters for each MTP block is only 0.33B.
@@
Table 1 summarizes detailed configurations of MiMo-V2-Flash. The model consists of 39 SWA layers and 9 GA layers. Both SWA and GA utilize Grouped-Query Attention (GQA) [ainslie2023gqa]. Specifically, SWA has 64 query heads and 8 key-value heads, while GA has 64 query heads and 4 key-value heads. The per-head dimensions are the same for SWA and GA (192 for queries and keys, and 128 for values). Rotary Positional Embedding (RoPE, su2024roformer) is partially applied to the first 64 dimensions query and key. Following recent best practices, we adopt an FP8 mixed-precision framework similar to DeepSeek-V3 [liu2024deepseek]. Specifically, we retain BF16 precision for the attention output projections, as well as for the embedding and output head parameters, while maintaining FP32 precision for the MoE router parameters. This mixed-precision configuration improves numerical stability without materially impacting training efficiency or memory footprint.
@@ 2.2 Hybrid Sliding Window Attention Architecture
Sliding window attention [beltagy2020longformer] restricts each token’s attention scope to a local window rather than the entire sequence, thereby reducing both computational and memory complexity dramatically. This naturally motivates hybrid attention architectures that interleave sliding window attention with global attention. However, prior work has shown that overly aggressive use of SWA, such as very small sliding window sizes or high SWA:GA ratios, can lead to substantial degradation in model performance [team2025gemma], especially in long-context tasks. Recently, the introduction of learnable attention sink bias, which allows the model to assign little or no attention to tokens when needed, has substantially enhanced the modeling capacity of SWA-based architectures [agarwal2025gpt]. While the precise theoretical underpinnings of the attention sink mechanism remain an active research area [xiao2023efficient, attnsink_study, qiu2025gatedattn, sun2024massive], we empirically observed that learnable attention sinks bias dramatically enhance the performance of hybrid SWA models, matching or even surpassing baselines with fully GA layers.
In MiMo-V2-Flash, our implementation follows the design used in gpt-oss [agarwal2025gpt], where a learnable attention sink bias $sink\in\mathbb{R}$ is applied to the denominator of softmax for each attention head. Specifically, let the attention logits between token $i$ and $j$ of one single head be:
@@ 2.2.1 Model Architecture Experiments
We evaluate four model architecture variants in a comparative setting. These include an all global attention (All GA) baseline, a hybrid SWA model with a 128-token window without attention sinks bias, and two hybrid SWA models augmented with attention sinks bias using window sizes of 128 and 512, respectively. All variants share the same training pipeline: pre-training on 250B tokens with an 8,192 sequence length, long context extension to 32,768 over an additional 40B tokens, followed by long-context SFT and reasoning SFT with chain-of-thought supervision. We evaluate model variants across benchmarks covering general capability, long-context understanding, and complex reasoning. General-domain results (Table 2) are obtained from pre-trained base models without long-context extension, evaluating general knowledge and reasoning on MMLU hendrycks2020measuring, BBH [suzgun2022challenging], TriviaQA [joshi2017triviaqa], GSM8K [cobbe2021training], MATH [hendrycks2021measuring], CMMLU [li2023cmmlu], and MBPP [austin2021program]. Long-context results (Table 3) evaluate long-context–extended base models on GSM-Infinite [zhou2025gsm], NoLiMa [modarressi2025nolima], and RULER-32k [hsieh2024ruler], and long-context SFT models on MRCR [vodrahalli2024michelangelo]. For GSM-Infinite and Nolima, we construct internal few-shot benchmarks to assess base models under controlled long-context settings. Complex reasoning results (Table 4) evaluate the reasoning SFT models on AIME24&25 [AIME], LiveCodeBench [jain2024livecodebench], and GPQA-Diamond [rein2024gpqa].
@@ 2.3.1 Motivation of using MTP
Prior work demonstrates that MTP is a powerful training objective that enhances training efficiency and model quality [gloeckle2024better, liu2024deepseek, xia2025mimo]. Beyond these training benefits, we place stronger emphasis on exploiting MTP as a native draft model for self-speculative decoding to deliver real-deployment speedup. In the following, we elaborate on how MTP accelerates inference from two perspectives: general LLM decoding speedup and RL training acceleration.
@@
MTP acceleration is particularly well-suited for RL training [miles2025], where the rollout phase consistently emerges as the dominant bottleneck due to the inference and decoding costs. MTP addresses two key challenges in RL training:
  • It enables efficient and effective RL with small batches. Current RL training relies on large-batch, off-policy algorithms to maximize throughput [schulman2017proximal, zheng2025group, liu2025deepseek]. However, on-policy training is generally more stable and effective, yet its small batches underutilize GPU resources. MTP mitigates this limitation by scaling token-level parallelism instead of batch size, making small-batch, on-policy RL training more practical.
  • It mitigates GPU idleness from long-tail stragglers. As the rollout phase progresses, long-tail stragglers that process long sequences with small batch sizes (often approaching 1) can cause significant GPU idleness [zhong2025optimizing, gao2025rollpacker]. In such scenarios, MTP enhances the computational efficiency of both attention and FFN, substantially reducing overall latency.
@@ 3 Pre-Training
The MiMo-V2-Flash pre-training corpus consists of 27 trillion tokens drawn from a diverse collection of high-quality sources, including public web content, books, academic papers, code, mathematics, and broader STEM materials. Our data processing pipeline largely follows that of MiMo-7B [xia2025mimo], with a deliberate shift toward data exhibiting long-range dependencies. In particular, we emphasize long-form web documents and carefully curated code corpora such as repository-level code, pull requests, issues, and commit histories to strengthen the model’s ability to capture extended contextual relationships and perform complex, multi-step reasoning.
@@ 3.3.1 Evaluation Setup
We evaluate MiMo-V2-Flash-Base on a series of benchmarks, encompassing various capabilities: (1) General language understanding and reasoning, including BBH [suzgun2022challenging], MMLU [hendrycks2020measuring], MMLU-Redux [gema2024we], MMLU-Pro [wang2024mmlu], DROP [dua2019drop], ARC [clark2018think], HellaSwag [zellers2019hellaswag], WinoGrande [sakaguchi2021winogrande], TriviaQA [joshi2017triviaqa], GPQA-Diamond [rein2024gpqa], SuperGPQA [du2025supergpqa], and SimpleQA [simpleqa]. (2) Mathematics reasoning. GSM8K [cobbe2021training], MATH [hendrycks2021measuring], and AIME [AIME] (2024 & 2025). (3) Coding. HumanEval+ [liu2023your], MBPP+ [liu2023your], CRUXEval [gu2024cruxeval], MultiPL-E [cassano2022multipl], BigCodeBench [zhuo2024bigcodebench], LiveCodeBench-v6 [jain2024livecodebench], and SWE-Bench [jimenez2023swe] (few-shot Agentless Repair [xia2024agentless]). (4) Chinese understanding. C-Eval [huang2023c], CMMLU [li2023cmmlu], and C-SimpleQA [he2025chinese]. (5) Multilingual understanding. GlobalMMLU [singh2025global], and INCLUDE [romanou2024include]. (6) Long context. NIAH-Multi [hsieh2024ruler], GSM-Infinite [zhou2025gsm] (5-shot, Hard Ops-{2,4,6,8,10}).
@@ 3.3.2 Evaluation Results
Table 5 presents a comprehensive comparison of MiMo-V2-Flash-Base against leading open-source base models [team2025kimik2, liu2024deepseek]. MiMo-V2-Flash-Base delivers competitive performance across most benchmarks and consistently outperforms peers on reasoning tasks (MMLU-Pro, GPQA-Diamond, AIME). On SWE-Bench, it even surpasses the substantially larger Kimi-K2-Base while using less than one-third the parameters, underscoring the strength of our approach for realistic code-agent tasks. However, constrained by its limited parameter count, we observe MiMo-V2-Flash exhibits lower knowledge capacity compared to larger models, as reflected in SimpleQA.
@@ 4.1 Multi-Teacher On-Policy Distillation (MOPD): A New Post-Training Paradigm
Rather than merging model parameters or generating static offline datasets from experts, we formulate multi-teacher knowledge integration as an on-policy reinforcement learning process. The student model samples from its own evolving distribution and receives token-level supervision from domain-specific teachers through KL divergence rewards [minillm, Agarwal2023OnPolicyDO, lu2025onpolicydistillation], effectively combining specialized capabilities without the traditional trade-offs (Table 7).
@@ 4.4 Technical Formulation of MOPD
By default, we combine the advantages of MOPD with other types of advantages, such as those computed using Outcome Reward Models (ORMs), including GRPO [shao2024deepseekmath]. Let $\hat{A}_{\text{ORM}}$ denote the advantages computed by the ORMs; the final advantages are given by:
@@ 4.5.1 Evaluation Setup
We evaluate MiMo-V2-Flash on MMLU-Pro [wang2024mmlu], GPQA-Diamond [rein2024gpqa], HLE Text-only [phan2025humanity], AIME 2025 [AIME], LiveCodeBench (2024.08-2025.04) [jain2024livecodebench], HMMT Feb. 2025 [balunovic2025matharena], Arena-Hard [arenahard2024], LongBench V2 [bai2025longbench], MRCR [vodrahalli2024michelangelo] ({2,4,8}-needles, maximum 128K), SWE-Bench Verified [jimenez2024swebench], SWE-Bench Multilingual [yang2025swesmith], Terminal-Bench, BrowseComp [wei2025browsecomp], $\tau^{2}$-Bench [barres2025tau].